OpenStudy (anonymous):

Claire has received scores of 85, 88, 87, and 85 on her algebra tests. What is the minimum score she must receive on the fifth test to have an overall test score average of at least 88? (Hint: The average of a list of numbers is their sum divided by the number of numbers in the list.)

OpenStudy (cwrw238):

let the required score be x , then 85 + 88 + 87 + 85 + x = 5 * 88 solve for x

OpenStudy (anonymous):

5*88 = 440 then divide 440 by 5 = 88 ? 88 my answer ?

OpenStudy (a_clan):

Formula for calculating average is given in the hint. Another useful information is that , in an equation Left hand Side(LHS) is always equal to Right Hand Side(RHS). So we assume : Let the 5th number be 'x'. Now we have to find the numerical value of x. According to the given formula: Average = (85 + 88 + 87 + 85 + x) / 5 Average is already given as 88. So, we substitute it in our equation. 88=(85 + 88 + 87 + 85 + x)/5 What remains now is simple calculation. This is how cwrw238 arrived at the above mentioned equation.

OpenStudy (anonymous):

so what is the correct answer ?

OpenStudy (a_clan):

88=(85 + 88 + 87 + 85 + x)/5 Multiply both sides by 5 88*5=(85 + 88 + 87 + 85 + x) 440 = (85 + 88 + 87 + 85 + x) 440 = 345 + x Subtract 345 from both LHS and RHS 95 = x This will be your answer.

OpenStudy (anonymous):

95 ?

OpenStudy (a_clan):

yes. It could not be explained any easier. Hope you understand the method as well.

OpenStudy (anonymous):

thank you and yes i do just making sure its correct answer.